OpenStudy (anonymous):

Which equations have the solution x = 6 (choose all that apply)?

OpenStudy (anonymous):

Which equations have the solution x = 6 (choose all that apply)? Answer x + 1 = 7 -x – 1 = -5 6(x + 2) = 2x – 5(x – 8) + 26 3x + 16 = 2x + 10